At the high velocity, drag force is proportional to the squared velocity of a particle as kv^2. Find its acceleration when a falling speed becomes half its terminal velocity. 



1
Expert's answer
2021-05-27T10:37:37-0400

Gives

Drag force

FD="\frac{1}{2}C\rho Av^2"

According to question

"F_D=kv^2"

Newton's 2ndlaw

F=Ma

Both force are equal

"Ma=kv^2"

"a=\frac{kv^2}{M}"

When

"v'=\frac{v}{2}"

"a'=?"

"a'=\frac{k(\frac{v}{2})^2}{M}=\frac{kv^2}{4M}=\frac{a}{4}"